Novak Incorporated leases a piece of equipment to Windsor Corporation on January 1, 2017. The lease agreement called for annual rental payments of $5,441 at the beginning of each year of the 4-year lease. The equipment has an economic useful life of 6 years, a fair value of $26,400, a book value of $21,400, and both parties expect a residual value of $8,100 at the end of the lease term, though this amount is not guaranteed. Novak set the lease payments with the intent of earning a 6% return, and Windsor is aware of this rate. There is no bargain purchase option, ownership of the lease does not transfer at the end of the lease term, and the asset is not of a specialized nature. Prepare the lease amortization schedule(s) for Windsor for all 4 years of the lease
Answer #1
Type of lease = Operating lease (Hint: Condition should not satisfied.)
bargain purchase option.
Ownership of the lease does transfer at the end of the lease term,.
Asset is a specialized nature.
Lease period is at least 75% of the useful life of the asset. (Lease period = 4/6 = 66.67%)
Present value of Minimum Lease payment is at least 90% of fair value of assets at the date of lease agreement. (19985/26400 = 75.70%)
Year PV factor @ 6%
0            1.00000
1            0.94340
2            0.89000
3            0.83962
Total            3.67301
Present value of lease liability = 5441* 3.67301) $          19,985
Windsor (Lessee)
Lease Amortization Schedule
Date Beginning Lease Liability Annual Lease Payment Plus BPO Interest on Liability Reduction of Lease Liability Lease Liability
January 1, 2017 $          19,985
January 1, 2017 $      19,985 $          5,441 $         5,441 $          14,544
January 1, 2018 $      14,544 $          5,441 $            873 $         4,568 $            9,976
January 1, 2019 $        9,976 $          5,441 $            599 $         4,842 $            5,133
January 1, 2020 $        5,133 $          5,441 $            308 $         5,133 $                     0
Total $        21,764 $        1,779 $       19,985
Interest on Liability = Beginning Lease Liability * 6%
Reduction of Lease Liability = Annual Lease Payment Plus BPO - Interest on Liability
Lease Liability = Beginning Lease Liability - Reduction of Lease Liability